2x/3 as a fraction minus 5= 5x/4 as a fraction +9
Solve for x.

Answer by Juhi(44)

2x/3 - 5 = 5x/4 + 9
2x/3 - 5x/4 = 9 + 5
-7x = 14
x = 14/-7
= -2